OpenStudy (anonymous):

Systems of Equations, Go! y = 3x − 4 y = −3x + 2

OpenStudy (anonymous):

first equation add the second equation 2y = -2 y = -1 -1 = 3x - 4 3 = 3x x = 1

OpenStudy (anonymous):

thank you